Aspirin is one of the products when you mix Salicylic acid and Acetic anhydride. One

kilogram of Salicylic acid produced one thousand and two hundred grams of Aspirin.

C7H6O3(s) + C4H6O3(l) C9H8O4(s) + CH3COOH(i)

a. Balance the chemical equation;

b. Identify the limiting reactant and the excess reactant if applicable;

c. Compute for the theoretical yield;

d. Determine the percent yield of the reaction;

e. Calculate the percent error and

f. Compute for the excess amount of the excess reactant if applicable.



1
Expert's answer
2021-11-03T04:19:03-0400

C7H6O3(S) + C4H6O3 → C9H8O4(S) + CH3CO2H

According to the reaction, n (C9H8O4) = n (C7H6O3)

n = m / M

M (C7H6O3) = 138.121 g/mol

M (C4H6O3) = 102.09 g/mol

M (C9H8O4) = 180.16 g/mol

n (C7H6O3) = 1 000 / 138.121 = 7.24 mol

n (C9H8O4)theoretical = n (C7H6O3) = 7.24 mol

m (C9H8O4)theoretical = n x M = 7.24 x 180.16 = 1304 g

%(C9H8O4) = 1200 / 1304 x 100 = 92%